# Moran Mizrahi1 Chen Shani2 Gabriel Stanovsky1 Dan Jurafsky2 Dafna Shahaf1 1

### v1

@@ Appendix C Text to Tree Parser Prompts
Given a recipe title, id, and ingredients, for each ingredient, determine: (1) Abbreviation: The shortest clear description. (2) Reference Type: Identify if the ingredient is for structure (‘structure’) or taste (‘taste’) of the dish. Ingredients impacting both are labeled as ’taste’. (3) Core Ingredient: Boolean indicating if the ingredient is essential to the identity of the dish (e.g., True for chocolate in chocolate cake). (4) Abstraction: Simplify the ingredient to its base form (e.g., ‘basil’ to ‘herb’, ‘walnuts’ to ‘nut’, ‘eggs’ to ‘egg’). Please return the results in the following JSON format only: {“recipe_id”: [(abbreviation, ref, core, abstraction), …], …}. INPUT: recipe_title, recipe_id, ingr_list, recipe_title, recipe_id, ingr_list, … OUTPUT:
@@ Appendix D Tree Edit Distance Implementation
To ensure the dishes are indeed recombinations, we discard any recipe that lacks at least one essential ingredient from both original dishes (if such an ingredient exists). We define an essential ingredient as one that appears frequently in recipes for a given dish but is not broadly common across all recipes (e.g., lasagna sheets in lasagna). Additionally, we remove ideas that are too similar to the seed recipes. To ensure that the combined tree preserves cross-dish inspiration, we require that at least 30% of its elements (nodes and edges) come from each original recipe.
@@ Appendix E Identifying Conflicting Ingredients
We remove ingredients that appear to collide in their flavors. Specifically, we look for pairs of ingredients that seldom appear together in the recipe repository, treating their pairing as uncommon and attempt to determine whether it might be a creative success or a failure.
To do so, we rely on two external datasets. First, we use flavorDB Garg et al. (2018), which catalogs taste molecules for a wide range of raw ingredients such as fruits, vegetables, and fish. Inspired by this dataset owner’s claim that two ingredients pair well if they share a larger proportion of taste molecules, we define a Jaccard-based pairing score between raw ingredients. Since flavorDB does not cover processed ingredients (e.g., lasagna sheets that consist of flour, water, and eggs), we also use FoodData Central Fukagawa et al. (2022) to infer their raw components. We define the pairing score between two composite ingredients as the low- est score among their constituent raw-ingredient pairs. After exploration, we chose 0.3 as our value threshold. If the score falls below 0.3, we consider the pairing problematic.
@@ Appendix F Tree to Text Prompts
This section includes the prompts we used to prompt GPT-4o to translate structured tree representation back into natural language. The system message for all prompts presented here was: “‘You are a cooking expert”.

### v2

@@ Appendix C Text to Tree Parser Prompts
[…]
@@ Appendix D Tree Edit Distance Implementation
To ensure the dishes are indeed recombinations, we discard any recipe that lacks at least one essential ingredient from both original dishes (if such an ingredient exists). We define an essential ingredient as one that appears frequently in recipes for a given dish but is not broadly common across all recipes (e.g., lasagna sheets in lasagna). Additionally, we remove ideas that are too similar to the seed recipes. To ensure that the combined tree preserves cross-dish inspiration, we require that at least 30% of its elements (nodes and edges) come from each of the original recipes.
@@ Appendix E Identifying Conflicting Ingredients
We remove ingredients that appear to collide in their flavors. Specifically, we look for pairs of ingredients that seldom appear together in the recipe repository, treating their pairing as uncommon and attempt to determine whether it might be a creative success or a failure. To do so, we rely on two external datasets. First, we use flavorDB Garg et al. (2018), which catalogs taste molecules for a wide range of raw ingredients such as fruits, vegetables, and fish. Inspired by this dataset owner’s claim that two ingredients pair well if they share a larger proportion of taste molecules, we define a Jaccard-based pairing score between raw ingredients. Since flavorDB does not cover processed ingredients (e.g., lasagna sheets that consist of flour, water, and eggs), we also use FoodData Central Fukagawa et al. (2022) to infer their raw components. We define the pairing score between two composite ingredients as the lowest score among their constituent raw-ingredient pairs. After exploration, we chose 0.3 as our value threshold. If the score falls below 0.3, we consider the pairing problematic.
@@ Appendix F Tree to Text Prompts
This section includes the prompts we used to prompt GPT-4o to translate structured tree representation back into natural language. The system message for all prompts presented here was: “You are a cooking expert”.